# Codes Correcting Transpositions of Consecutive Symbols

### v1

@@ 5 Codes correcting all patterns of transpositions
Figure 1: Lower bounds on the asymptotic rate of optimal transposition-correcting codes of relative distance $\delta$ over an alphabet of size $q=4$; see (70). The bound from Theorem 5.1 is better than that from Theorem 4.1 for $\delta>\delta_{0}\approx 0.343$.
@@ 6 Conclusion and further work
We have analyzed the problem of correcting transpositions of consecutive symbols in $q$-ary strings—types of errors that may occur in various contexts, e.g., as spelling errors, timing errors, inter-symbol interference errors (such as bit-shifts in magnetic storage devices), etc. As one of the main questions left open in this line of work we mention a construction of codes correcting $t>1$ transpositions that are asymptotically optimal in the sense that their cardinality scales as $\Theta(\frac{q^{n}}{n^{t}})$ when $n\to\infty$ (see (39)). In the linear regime $t=\tau n$, we have presented a lower bound on the achievable asymptotic code rates. Since the code space is nonuniform, obtaining good upper bounds is a nontrivial problem that is also left for future work, as are explicit constructions of asymptotically good (i.e., having positive rate) codes correcting a fraction of $\tau$ transpositions. For codes of infinite minimum distance, the problem of deriving the exact value of the maximum achievable rate (i.e., the zero-error capacity) is very interesting, and is still open for all alphabet sizes.
Finally, as mentioned in Section 1.3, there are several related models describing communication in the presence of synchronization and reordering errors that are potentially relevant for modern information transmission and storage systems and that are not yet fully understood, e.g., timing channels, non-binary channels with successive transpositions, permutation channels with bounded displacements, channels with a combination of transpositions and other types of errors, etc.

### v2

@@ 3.2 Codes correcting an arbitrary number of transpositions and the zero-error bound
Figure 1: Lower bounds on the asymptotic rate of optimal codes correcting a fraction of $\tau$ transpositions over an alphabet of size $q=4$; see (45). The bound from Theorem 3.2 is better than that from Theorem 3.1 for $\tau>\tau_{0}\approx 0.172$.
4 Bounds on codes correcting $t=\textnormal{const}$ transpositions
In this section we present bounds on the cardinality of optimal codes in $\mathbb{Z}_{q}^{n}$ correcting $t=\textnormal{const}$ transpositions (Theorem 4.1 ahead). In contrast with the linear $t=\tau n$ regime, the lower bound obtained via the Gilbert–Varshamov argument is not better than the bounds that can be obtained by different methods, see Remarks 3 and 4 ahead. Nonetheless, we derive the expression for the average ball size and state the corresponding bound in the $t=\textnormal{const}$ regime as they may be of separate interest.
Lemma 1
For fixed $\mu\in[0,1]$,
$\lim_{n\to\infty}\frac{1}{n}\log_{q}\!\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\colon\operatorname{run}(\boldsymbol{x})\leqslant\mu n\}\right|=\begin{cases}H_{q}(\mu),&0\leqslant\mu<1-q^{-1}\\ 1,&1-q^{-1}\leqslant\mu\leqslant 1\end{cases},$ (46)
where $H_{q}(\cdot)$ is the $q$-ary entropy function.
Proof
Let $S_{q}(n,m)$ denote the number of strings from $\mathbb{Z}_{q}^{n}$ having exactly $m$ runs of identical symbols, so that
$\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\colon\operatorname{run}(\boldsymbol{x})\leqslant\mu n\}\right|=\sum_{m=1}^{\lfloor\mu n\rfloor}S_{q}(n,m).$ (47)
By a direct counting argument we find that
$S_{q}(n,m)=\binom{n-1}{m-1}q(q-1)^{m-1}$ (48)
and hence, by Stirling’s approximation,
$\lim_{n\to\infty}\frac{1}{n}\log_{q}S_{q}(n,\lfloor\mu n\rfloor)=H_{q}(\mu).$ (49)
$\max_{m\leqslant\lfloor\mu n\rfloor}S_{q}(n,m)\;\leqslant\;\sum_{m=1}^{\lfloor\mu n\rfloor}S_{q}(n,m)\;\leqslant\;\mu n\cdot\max_{m\leqslant\lfloor\mu n\rfloor}S_{q}(n,m),$ (50)
its exponential growth-rate (to base $q$) is
$\lim_{n\to\infty}\frac{1}{n}\log_{q}\sum_{m=1}^{\lfloor\mu n\rfloor}S_{q}(n,m)=\max_{0\leqslant\mu^{\prime}\leqslant\mu}H_{q}(\mu^{\prime}).$ (51)
That this is equal to the right-hand side of (46) follows from the fact that $H_{q}(\mu)$ is a concave function of $\mu\in[0,1]$ maximized at $\mu=1-q^{-1}$, the maximum being $H_{q}(1-q^{-1})=1$. $\blacksquare$
where $H(1-q^{-1}-\epsilon)<1$. A very useful consequence of this fact is that one can disregard such strings in the asymptotic analysis of optimal codes in the $n\to\infty$, $t=\textnormal{const}$ regime. Namely, since $M^{\star}_{q}(n;t)$ is lower-bounded by the cardinality of optimal codes correcting $t$ substitutions (see Section 1.3), and the latter for constant $t$ scales as $\Omega(\frac{q^{n}}{n^{2t}})$ (by the standard Gilbert–Varshamov bound in the Hamming metric [15, Section 2.8]), we conclude that
$M^{\star}_{q}(n;t)=\Theta\!\left(q^{n+o(n)}\right).$ (53)
Relations (52) and (53) imply that, for an optimal code correcting $t$ transpositions, throwing out its codewords that have at most $(1-q^{-1}-\epsilon)n$ runs affects its cardinality by an asymptotically negligible amount.
Theorem 4.1
For any fixed $q\geqslant 2$ and $t\geqslant 1$, as $n\to\infty$,
$\frac{(2t)!\,q^{2t}}{(q-1)^{2t}}\frac{q^{n}}{n^{2t}}\;\lesssim\;M^{\star}_{q}(n;t)\;\lesssim\;\frac{t!\,q^{t}}{(q-1)^{t}}\frac{q^{n}}{n^{t}}.$ (54)
Proof
From Proposition 6 we can find the asymptotic behavior of $T^{\textnormal{e}}_{q}(n;r)$ for fixed $r$ and $n\to\infty$,
$T^{\textnormal{e}}_{q}(n;r)\;\sim\;n^{r}\frac{1}{r!}q^{n-r}(q-1)^{r},$ (55)
where we used the fact that $\binom{n}{s}\sim\frac{n^{s}}{s!}$ when $s$ is fixed and $n\to\infty$. From (55), (17) and (18) it follows that $T_{q}(n;r)$ and $\bar{T}_{q}(n;r)$ have the same asymptotic scaling as $T^{\textnormal{e}}_{q}(n;r)$ in the regime currently of interest. This fact, together with the bound ${M}_{q}^{\star}(n;t)\geqslant\frac{q^{2n}}{\bar{T}_{q}(n;2t)}$ (see (31)) implies the left-hand side of (54).
In order to show the right-hand inequality, an asymptotic lower bound on the cardinality of $B(\boldsymbol{x};t)$ (the set of all strings that can be obtained by applying at most $t$ transpositions to the string $\boldsymbol{x}$) will be needed. Using Proposition 4, for strings satisfying $\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n$, $|B(\boldsymbol{x};t)|$ can be asymptotically lower-bounded by
$\gtrsim\;\sum_{u=0}^{t}\binom{(1-q^{-1}-\epsilon)n/2}{u}\binom{(1-q^{-1}-\epsilon)n/2}{t-u}$ (56)
$\sim\;\sum_{u=0}^{t}\frac{(1-q^{-1}-\epsilon)^{u}\,n^{u}}{2^{u}\,u!}\frac{(1-q^{-1}-\epsilon)^{t-u}\,n^{t-u}}{2^{t-u}\,(t-u)!}$
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{2^{t}}\,\sum_{u=0}^{t}\frac{1}{u!\;(t-u)!}$
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{2^{t}\,t!}\,\sum_{u=0}^{t}\binom{t}{u}$
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{t!}.$
We can now derive the upper bound in (54) by a packing argument. Let $\mathcal{C}^{\star}_{q}(n;t)$ be an optimal code correcting $t$ transpositions, meaning that $|\mathcal{C}^{\star}_{q}(n;t)|=M^{\star}_{q}(n;t)$. By (3), the sets $B(\boldsymbol{x};t)$ centered at codewords of $\mathcal{C}^{\star}_{q}(n;t)$ do not overlap, so we have
$\sum_{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n;t)\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n}|B(\boldsymbol{x};t)|\;\leqslant\;\sum_{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n;t)}|B(\boldsymbol{x};t)|\;\leqslant\;|\mathbb{Z}_{q}^{n}|\;=\;q^{n}.$ (57)
Together with (56) this implies that
$\frac{1}{t!}\left(1-q^{-1}-\epsilon\right)^{\!t}n^{t}\cdot\left|\left\{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n;t)\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n\right\}\right|\;\lesssim\;q^{n},$ (58)
From Lemma 1 and the discussion thereafter we know that we can safely ignore in the asymptotic analysis the codewords with at most $(1-q^{-1}-\epsilon)n$ runs, meaning that
$\left|\left\{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n;t)\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n\right\}\right|\;\sim\;\left|\mathcal{C}^{\star}_{q}(n;t)\right|\;=\;M^{\star}_{q}(n;t).$ (59)
Now (58) and (59) imply
$\frac{1}{t!}\left(1-q^{-1}-\epsilon\right)^{\!t}n^{t}\cdot M^{\star}_{q}(n;t)\;\lesssim\;q^{n}.$ (60)
Since (60) holds for any $\epsilon>0$, the upper bound in (54) is established. $\blacksquare$
Remark 3(Lower bound for the binary case)
For $q=2$, a lower bound better than the one in Theorem 4.1 was derived in [21, Theorem 11] by using an interpretation of binary codes correcting successive transposition errors as packings in the Manhattan metric, and using the densest known packings, namely
$M^{\star}_{2}(n;t)\;\gtrsim\;c_{t}\frac{2^{n+t}}{n^{t}},$ (61)
where $c_{1}=\frac{1}{2}$, $c_{2}=\frac{1}{3}$,333[21, Theorem 11] states a slightly worse constant $c_{2}=\frac{1}{4}$. It can be shown based on [30, Theorem 3] that $c_{2}$ can be improved to $\frac{1}{3}$. and $c_{t}=\frac{1}{2t+1}$ for $t\geqslant 3$. In particular, in the binary case we have
$M^{\star}_{2}(n;t)=\Theta\Big(\frac{2^{n}}{n^{t}}\Big).$ (62)
Unfortunately, as already mentioned in Remark 2, this geometric interpretation fails when $q>2$ and thus the construction that implies this lower bound does not generalize to larger alphabets. $\blacklozenge$
Remark 4(Lower bound via substitution-correcting codes)
For general $q$, a lower bound better than the one in (54) follows from the known bounds for substitution-correcting codes (see Section 1.3). For example, it is known that $q$-ary codes correcting a single substitution and having cardinality $\Theta(\frac{q^{n}}{n})$ exist [15] (e.g., Hamming codes), and hence
$M^{\star}_{q}(n;1)\;=\;\Theta\Big(\frac{q^{n}}{n}\Big).$ (63)
For $t>1$, Varshamov’s improvement of Gilbert’s bound [15, Section 2.9] implies that
$M^{\star}_{q}(n;t)\;=\;\Omega\Big(\frac{q^{n}}{n^{2t-1}}\Big).$ (64)
Further improvements of the exponent in the denominator are possible in some cases; see, e.g., [8, 34].
Based on what is known in the special case $t=1$ (see (63)), as well as in the case of general $t$ and $q=2$ (see (62)), it is reasonable to conjecture that
i.e., that the redundancy of $t\log_{q}n+{\mathcal{O}}(1)$ is achievable for any fixed $t$ and alphabet size $q$. At present, however, we do not have a proof of this fact. $\blacklozenge$
@@ 5 Conclusion and further work
We have analyzed the problem of correcting transpositions of consecutive symbols in $q$-ary strings—types of errors that may occur in various contexts, e.g., as synchronization and timing errors, spelling errors, inter-symbol interference errors, genome rearrangement errors, etc.
In the linear regime $t=\tau n$, we have presented a Gilbert–Varshamov-like lower bound on the achievable code rates. Since the code space is nonuniform, obtaining good upper bounds is a nontrivial problem that is left for future work, as are explicit constructions of asymptotically good (i.e., having positive rate) codes correcting a fraction of $\tau$ transpositions. For codes of infinite minimum distance, the problem of deriving the exact value of the maximum achievable rate, i.e., the zero-error capacity, is very interesting and is still open for all alphabet sizes. In the $t=\textnormal{const}$ regime, as one of the main open questions in this line of work we mention a construction of nonbinary codes correcting $t>1$ transpositions whose cardinality scales as $\Theta(\frac{q^{n}}{n^{t}})$ when $n\to\infty$ (see (65)).
Finally, there are several related models describing communication in the presence of synchronization and reordering errors that are potentially relevant for modern information transmission and storage systems and that are not yet fully understood, e.g., timing channels, permutation channels with bounded displacements, channels with a combination of transpositions and other types of errors, etc.